# Risk-Neutral Pricing Model of Uniswap Liquidity Providing Position: A Stopping Time Approach

### v1

@@ III-C European LP Pricing
$=\mathbb{E}[e^{-r\tau}\text{LP}_{H}|S(\tau)=S_{H}]+\mathbb{E}[e^{-r\tau}\text{LP}_{L}|S(\tau)=S_{L}]$
$\mathbb{+}E(V_{\text{fee}}(t))$
$=\text{LP}_{H}\mathbb{E}[e^{-r\tau}|S(\tau)=S_{H}]+\text{LP}_{L}\mathbb{E}[e^{-r\tau}|S(\tau)=S_{L}]$
@@ III-E Rebate Pricing
[…]
$=C_{a}L_{q}\cdot\mathbb{E}\left[\frac{1-e^{-r\tau}}{r}\right]$
$=\frac{C_{a}L_{q}}{r}\cdot(1-\mathbb{E}\left[{e^{-r\tau}}\right])$
@@
$\mathbb{E}(V_{\text{fee}}(t))=\mathbb{E}(C_{a}L_{q}e^{-r\tau}\tau)=C_{a}L_{q}\mathbb{E}(e^{-r\tau}\tau)$
@@ IV Sensitivity Analysis
Figure 2: Table I market condition pv plot
Figure 3: Table II market condition pv plot
IV-A European and American Present Value Comparison
We sampled two sets of market parameters based on actual market conditions to observe how the European and American pricing models reflect the actual value of Uniswap range order positions. As illustrated in Table II, Through numerical analysis, it can be observed that under the market parameters (C=0.04, r=0.05, $\sigma$=0.4), if we immediately establish a Uniswap range order position and exit the position only when it touches the boundary (i.e., behavior under European LP pricing), the present value of such a product will certainly be less than 1. In contrast, if we are able to exit at the optimal execution boundary before reaching the designated boundary, even without achieving a high value of C, we may still obtain a positive return. This is because the American LP pricing model indicates that if a better execution exit boundary exists outside the set LH, a positive return expectation is possible. Furthermore, when C increases (e.g., C=0.2, r=0.05, $\sigma$=0.7), As illustrated in Table I, even with a high value of $\sigma$ (which implies that we may reach the boundary and exit the position more quickly), the European LP pricing model still yields a positive expected value under the current spot price of 1. We also plotted the PV-spot graphs for both European and American pricing models under two sets of market parameters. Figures 2 and Figures 3 clearly demonstrate that, although the early exit boundary for the American option pricing model is narrower, it yields a significantly higher positive expected return compared to the European pricing model.
Figure 4: delta comparison between european and payoff
IV-B Delta Analysis
Delta refers to the sensitivity of the option’s price with respect to changes in the underlying asset’s price. As illustrated in Figure 4, the delta for European pricing and payoff valuation reveals noteworthy disparities due to inherent market parameters (C=0.05, r=0.04, $\sigma$=0.25). At the orange line’s zero point, corresponding to the range (1, 1.05) in the previous analysis, the delta suggests that during actual trading, the delta derived from payoff valuation might lead to residual risks characterized by the Greek value. This indicates potential risk exposure when delta hedging is performed under the assumption of risk neutrality.
IV-C The Impact of Volatility

### v2

@@ III-C European LP Pricing
$=\mathbb{E}[e^{-r\tau}V_{\text{LP}}(H)|S(\tau)=S_{H}]+\mathbb{E}[e^{-r\tau}V_{\text{LP}}(L)|S(\tau)=S_{L}]$
$+\mathbb{E}(V_{\text{fee}}(t))$
$=V_{\text{LP}}(H)\mathbb{E}[e^{-r\tau}|S(\tau)=S_{H}]+V_{\text{LP}}(L)\mathbb{E}[e^{-r\tau}|S(\tau)=S_{L}]$
@@ III-E Rebate Pricing
$=C_{a}L_{q}V_{LP}(P_{t})(\cdot\mathbb{E}\left[\int_{0}^{\tau}e^{-rt}dt\right]$
$=C_{a}L_{q}V_{LP}(P_{t})\cdot\mathbb{E}\left[\frac{1-e^{-r\tau}}{r}\right]$
$=\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}\cdot(1-\mathbb{E}\left[{e^{-r\tau}}\right])$
@@
$\mathbb{E}(V_{\text{fee}}(t))=\mathbb{E}(C_{a}L_{q}V_{LP}(P_{t})e^{-r\tau}\tau)=C_{a}L_{q}V_{LP}(P_{t})\mathbb{E}(e^{-r\tau}\tau)$
@@
Figure 2: Euro-Uniswap price distribution: Monte Carlo vs analytical solution
@@
Figure 2 provides a Monte Carlo simulation to model the price dynamics of an asset following a Geometric Brownian Motion (GBM). The initial asset price $S_{0}$ was set to 1, with volatility $\sigma=0.6$, drift $\mu=0$, risk-free rate $r=0.04$, and boundary levels $H=1.2$ (upper bound) and $L=0.8$ (lower bound). We ran the simulation 10,000 times, calculating the payoff at each simulation based on whether the price reached the upper or lower boundary, then discounting the payoff to obtain the expected value. The Monte Carlo estimates closely matched the analytical solution derived from our new pricing formula, confirming the accuracy of our approach.
The final distribution of the outcomes, represented by a histogram, exhibited a bimodal shape. This reflects the nature of the problem, where the asset price is either stopped at the upper bound $H$ or the lower bound $L$ due to the stopping conditions inherent in the Uniswap V3 liquidity model. The two peaks correspond to the outcomes where the price reaches these boundaries, capturing the distinctive characteristics of liquidity provision in a decentralized exchange framework. This bimodal distribution highlights the risk/reward trade-offs faced by liquidity providers in such a system, where extreme price movements lead to clear-cut outcomes based on boundary interactions.
III-F Formulation of Case-Specific Models
Based on the derivations from the foundational section, we have obtained the pricing formulas for the following four scenarios. For the pricing formula of the European LP option problem under a GBM process, assuming that transaction fees can only be collected upon exiting the position, the pricing formula is as follows:
$V_{t}$
$=LP_{H}e^{\mu^{\prime}b^{\prime}}\frac{\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L}e^{\mu^{\prime}a^{\prime}}\frac{\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+C_{a}L_{q}V_{LP}(P_{t})$
$(\frac{(b-a)(e^{\mu^{\prime}a^{\prime}}\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}}))}{\sqrt{2r+\mu^{\prime 2}}\operatorname{th}((b-a)\sqrt{2r+\mu^{\prime 2}})\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$-\frac{e^{\mu^{\prime}a^{\prime}}b^{\prime}\operatorname{ch}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}a^{\prime}\operatorname{ch}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\sqrt{2r+\mu^{\prime 2}}\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})})$
When we assume that transaction fees can be continuously collected during the lifetime of the position, the pricing formula becomes:
$V_{t}$ $=LP_{H}e^{\mu^{\prime}b^{\prime}}\frac{\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L}e^{\mu^{\prime}a^{\prime}}\frac{\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}(1$
$-\frac{e^{\mu^{\prime}a^{\prime}}\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})})$
For the pricing formula of the American LP option problem, assuming that transaction fees can only be collected upon exiting the position, the pricing formula is as follows:
$V_{t}=LP_{L_{2}}e^{\mu^{\prime}d^{\prime}}\frac{\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L_{1}}e^{\mu^{\prime}c^{\prime}}\frac{\operatorname{sh}(bd^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}+C_{a}L_{q}V_{LP}(P_{t})($
$\frac{(d-c)(e^{\mu^{\prime}c^{\prime}}\operatorname{sh}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}}))}{\sqrt{2r+\mu^{\prime 2}}\operatorname{th}((d-c)\sqrt{2r+\mu^{\prime 2}})\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$-\frac{e^{\mu^{\prime}c^{\prime}}d^{\prime}\operatorname{ch}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}c^{\prime}\operatorname{ch}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\sqrt{2r+\mu^{\prime 2}}\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})})$
When we assume that transaction fees can be continuously collected during the lifetime of the position, the pricing formula for the American LP option problem becomes:
$V_{t}$ $=LP_{L_{2}}e^{\mu^{\prime}d^{\prime}}\frac{\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L_{1}}e^{\mu^{\prime}c^{\prime}}\frac{\operatorname{sh}(bd^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$+\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}(1$
$-\frac{e^{\mu^{\prime}c^{\prime}}\operatorname{sh}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})})$
@@ IV Sensitivity Analysis
Figure 3: European LP Pricing Analysis
Our subsequent Greeks analysis will be conducted based on two sets of market parameters corresponding to different market conditions. One scenario features high transaction fees and high market volatility, referred to as ’good market situation’, as these conditions allow Uniswap V3 positions to accrue more fees. The other scenario involves low transaction fees and low market volatility, termed ’good market situation’.
Taking the European LP option pricing model as an example, we present the valuation discrepancies between pricing under good/bad market conditions and the payoff model at different price levels. The good/bad market conditions referenced here are based on the market parameters used in Table I and Table I. The range order price interval is set at (0.8,1.2). As shown in Figure 3, under the calculations based on the European LP option pricing model, there is a significant difference between the current Uniswap V3 payoff value and its actual present value. This is because the current value of a Uniswap V3 position is not limited to the extractable payoff; it also incorporates the discounted value of future fees that can still be collected if the position is not immediately withdrawn.
In scenarios with relatively low volatility and low transaction fees, the analysis shows that if the Uniswap V3 position is not withdrawn now, its future value will depreciate, and the discounted value of future fees will be insufficient to offset the loss caused by the position’s depreciation, as depicted by the green line in the Figure 3. Conversely, under good market conditions with higher volatility and transaction fees, the orange line in the Figure 3 demonstrates that the actual value of the Uniswap V3 position exceeds its payoff value.
Figure 4: European LP Delta Analysis
IV-A Delta Analysis
In historical Uniswap V3 trading practices, to mitigate the uncertainty of future returns, delta hedging is often employed to maintain delta-neutral risk exposure for Uniswap V3 positions. However, based on our European LP option pricing model, we find that delta risk within the $(L,H)$ range differs from the delta derived from payoff-based calculations.
Taking the comparison between the orange and blue lines in Figure 4 as an example (since it is uncommon to persist with Uniswap V3 positions under unfavorable market conditions), we observe that when the price is at 0.9, the delta value of the orange line exceeds that of the blue line. This indicates that the actual delta risk at this point is higher than the delta risk derived from payoff-based pricing. Consequently, more delta needs to be hedged to achieve delta neutrality. Conversely, when the price is at 1.1, the actual delta risk associated with holding the Uniswap V3 position is lower. This suggests that hedging excessive delta based on payoff-based calculations is unnecessary.
Figure 5: European LP Gamma Analysis
IV-B Gamma Analysis
By conducting a gamma analysis of the European LP pricing model versus the payoff-based pricing model, we can observe from Figure 5 that holding a Uniswap V3 position inherently involves gamma risk. Gamma risk refers to the situation where large gamma values cause significant fluctuations in delta with small changes in the underlying price $S$. In such scenarios, the delta hedging strategy becomes less effective, as even minor changes in $S$ can lead to substantial delta risk, necessitating more frequent adjustments to maintain delta neutrality. Conversely, when gamma values are small, delta remains relatively stable, reducing the need for frequent hedging.
As shown in the Figure 5, there is a significant difference between the gamma risk values derived from payoff-based modeling and those from the European LP pricing model. This highlights the higher requirements for managing gamma risk in Uniswap V3 positions, emphasizing the need for more refined risk control strategies.
IV-C Vega Analysis
Vega measures the sensitivity of the valuation of a Uniswap V3 European pricing position to changes in volatility. From the Figure 6, we observe the variation of vega with respect to spot prices under different market conditions. Notably, there is no vega estimation graph for the payoff-based pricing model, as this model assumes no correlation between price and volatility, and thus does not account for vega risk.
From the Figure 6, it is evident that vega remains consistently negative within the defined $(L,H)$ range. This indicates that holding a Uniswap V3 position inherently represents a short volatility strategy, implying that lower volatility leads to higher returns.
Figure 6: European LP Vega Analysis
IV-D Greeks Comparison between different range order
After analyzing the unique patterns exhibited by the Greeks under Uniswap V3 positions, we aim to compare the Greeks of Uniswap V3 positions established with different $(L,H)$ ranges to uncover the relationship between Uniswap V3 pricing and the choice of $(L,H)$. The results are illustrated in the figure.
We set up three groups of Uniswap V3 range orders with price ranges of $(0.8,1.1)$, $(0.8,1.2)$, and $(0.8,1.3)$. The assumed market conditions are $r=0.05$, $C=0.2$, $\mu=0$, and $\sigma=0.7$. From Figure 7 ,we observe that as the pricing range widens, the European LP pricing model’s valuation results tend to increase. This is because a wider pricing range implies that the asset is expected to remain within the pricing range for a longer period, thereby extending the expected time for fee collection.
Figure 7: Different Range LP Pricing Comparison
An interesting observation arises from the delta comparison in Figure 8 . We find that as the pricing range widens, the delta values generally increase. This is because when the pricing range is very wide, the proportional impact of price fluctuations on the exchange between the underlying asset and currency diminishes relative to the total position value. Consequently, the position value remains strongly correlated with the underlying asset price, resulting in a higher delta. Conversely, when the pricing range is narrow, even small price movements can trigger significant exchange activity, weakening the correlation between position value and underlying asset price, thus reducing delta.
Figure 8: Different Range LP Pricing delta Comparison
Although wider pricing ranges lead to larger delta risks, the absolute value of gamma decreases, as shown in Figure 9. This indicates that delta hedging becomes easier for wider pricing ranges compared to narrower ones. With a broader pricing range, delta changes more smoothly as the spot price fluctuates, simplifying the delta hedging process.
Figure 9: Different Range LP Pricing gamma Comparison
Finally, as demonstrated in Figure 10, even with changes to the pricing range, vega remains strictly negative. This confirms that the relationship between price and volatility consistently exhibits a negative correlation.
Figure 10: Different Range LP Pricing Pricing Analyiss versus vol
To validate the effectiveness of our Greeks, we need to perform numerical tests. Specifically, based on the Taylor expansion method, the following equation holds:
$\Delta V\approx\text{Delta}\times\Delta S+\frac{1}{2}\text{Gamma}\times\Delta S^{2}+\text{Vega}\times\Delta\sigma$
We can use two sets of different parameters, $S$ and $\sigma$, to calculate the valuation change of our Uniswap V3 position before and after market fluctuations (based on the present value derived from our pricing model). If the Greeks we calculate, when integrated into the right-hand side of the equation, can estimate the pricing with an acceptable level of accuracy, it will sufficiently demonstrate the validity of our Greeks.
The parameters we set are as follows, for the market before the fluctuation:
$C=0.2,r=0.05,\mu=0,\sigma=0.7$
$S=1,H=1.1,L=0.8$
Taking ETH as an example, under general circumstances, we assume a 5 % decrease in price and a 10 % increase in volatility itself. Using our pricing model, the valuation of the Uniswap V3 position before and after the market fluctuation is calculated as follows:
The Current Pv is: 1.02110809997664861.0211080999766486 1.02110809997664861.0211080999766486
The new PV is: 0.99304363754849550.9930436375484955 0.99304363754849550.9930436375484955
The Position Changing is $-0.028064462428153192$
The Greeks obtained through calculation, along with the corresponding Taylor expansion results, are as follows:
The Current delta is: 0.197906601375752440.19790660137575244 0.197906601375752440.19790660137575244
The Current gamma is: $-11.101273578173476$
The Current vega is: $-0.15958993211606654$
Taylor Expansion estimation is: $-0.034943217289629125$
The Current residential is: $-0.006878754861475933$
Thus, under this level of market changes, the pricing estimation derived from the Taylor expansion exhibits an approximate error of 0.6% compared to the actual valuation of the Uniswap V3 position.
IV-E The Impact of Volatility